# First Measurement of Charged Current Muon Neutrino-Induced $K^{+}$ Production on Argon using the MicroBooNE Detector

### v3

@@ Abstract
Within the Standard Model (SM), the stability of protons is a consequence of the global symmetry that conserves baryon number. However, various Grand Unified Theories (GUTs) predicting baryon number violation through nucleon decay suggest different decay channels, with a preference for $p\rightarrow e^{+}\pi^{0}$, with lifetimes ranging from $10^{34}$ to $10^{39}$ years Georgi and Glashow (1974). Similarly, the Supersymmetric Grand Unified Theory (SUSY GUT) favors protons decaying via $p\rightarrow\nu K^{+}$, with lifetimes of $10^{32}$ to $10^{35}$ years Hisano et al. (1993).
Upcoming large long-baseline neutrino oscillation experiments, such as the Deep Underground Neutrino Experiment (DUNE) Abi et al. (2020) and Hyper-Kamiokande Abe et al. (2018), have the sensitivity to search for proton decay in the range of $10^{32}$ to $10^{39}$ years. Cherenkov detector experiments like Super-Kamiokande Abe et al. (2014, 2017, 2020) and Hyper-Kamiokande are particularly suited for $p\rightarrow e^{+}\pi^{0}$ decay searches. However, for water Cherenkov detectors, the $p\rightarrow\nu K^{+}$ decay channel needs to be indirectly detected since the momentum of the produced $K^{+}$ (339 MeV/c) is below the Cherenkov detection threshold in water (560 MeV/c). Furthermore, backgrounds from atmospheric neutrino interactions, such as the $\nu_{\mu}n\rightarrow\mu^{-}K^{+}n$ and $\nu p\rightarrow\nu K^{+}\Lambda$ channels, are challenging to distinguish in water Cherenkov detectors.
@@
The first measurements of charged-current (CC) and neutral-current (NC) kaon production, reported with the Argonne National Laboratory Barish et al. (1974) and GARGAMELLE bubble chambers Deden et al. (1975) in 1974 and 1975, observed $\mathcal{O}$(10) events. Later, in the 1980’s, the Brookhaven National Laboratory bubble chamber reported $\mathcal{O}$(10) CC and NC events with strange particle production in neutrino-hydrogen and neutrino-deuterium interactions Jensen (1964); Baker et al. (1981), and the Fermilab bubble chamber reported $\mathcal{O}$(10) events and a measured cross section for CC neutrino interactions $\nu_{\mu}+n\xrightarrow{}\mu^{-}+\Lambda+K^{+}$ on a deuterium target for energy ranges between 10 GeV and 250 GeV Son et al. (1983). Most recently, the MINERvA experiment measured $K^{+}$ production in NC and CC modes on a hydrocarbon target, reporting $\mathcal{O}$(100) and $\mathcal{O}$(1000) events respectively Marshall et al. (2016, 2017); Marshall (2016). In addition, MINERvA reported the first evidence of coherent kaon production Wang et al. (2016); Marshall (2016).
@@
The MicroBooNE detector Fleming (2012); Acciarri et al. (2017) consists of an 85 tonne LArTPC and a photon detection system comprised of 32 photomultiplier tubes (PMTs). The detector identifies neutrino interactions by detecting ionization and scintillation light produced by charged particles moving through the detector. Ionization charge is recorded across three wire planes (two induction, one collection) oriented at different angles (with induction planes at $\pm 60^{\circ}$ relative to the vertical collection plane), and 3 mm wire-spacing. This configuration allows the generation of mm-scale resolution, three-dimensional images of neutrino interactions. Scintillation light collected by the PMTs provides ns-scale timing information, essential for identifying neutrino interactions that are synchronized with the BNB, and for rejecting cosmic-ray backgrounds.
@@
In MicroBooNE, the LArSoft framework Snider and Petrillo (2017) provides tools for simulating the neutrino flux, neutrino interactions, particle propagation, and detector response including ionization and scintillation processes as well as reconstruction of neutrino events. The BNB neutrino flux through the MicroBooNE detector is simulated using the MiniBooNE flux model Aguilar-Arevalo et al. (2009) adapted for MicroBooNE’s location along the beamline. GENIE v3.0.6 (G18_10a_02_11a) Andreopoulos et al. (2010, 2021) is the event generator used to simulate neutrino-argon interactions, including both single and associated kaon production. In GENIE, associated kaons are produced via individual resonances simulated by the Rein-Sehgal model Rein and Sehgal (1981). Kaons produced via deep inelastic scattering are simulated by the Koba-Nielsen-Olesen (KNO) scaling parameterization Koba et al. (1972). For higher energies (hadronic invariant mass W >> 2.3 GeV), the AGKY model Yang et al. (2009) and PYTHIA6 Sjöstrand et al. (2006) are used. The hadronization models are tuned to match strange particle production ($\Lambda$ and $K^{0}_{S}$) observed in experiments such as the Big European Bubble Chamber Jones et al. (1993); Allasia et al. (1983); Willocq et al. (1992); Bosetti et al. (1982) and the Fermilab bubble chamber Baker et al. (1986); DeProspo et al. (1994). The single kaon production model Rafi Alam et al. (2010) is implemented in GENIE, generating events in the channels $\nu_{l}+n\rightarrow l^{-}+n+K^{+}$, $\nu_{l}+p\rightarrow l^{-}+p+K^{+}$, and $\nu_{l}+n\rightarrow l^{-}+p+K^{0}$. The region of validity assumed for the model is for $E_{\nu}\leq$ 2 GeV Rafi Alam et al. (2010); Alvarez-Ruso et al. (2013), while the GENIE authors note that the model can be applied to higher energies as well Alam et al. (2015).
@@
Particle trajectories and interactions within the detector are simulated using GEANT4 v4_10_3_p03c using the QGSP_BERT hadron physics list Agostinelli et al. (2003); Allison et al. (2006, 2016), which models particle propagation in the detector. Simulated neutrino interactions are overlayed with data events collected using an unbiased trigger that operates in anti-coincidence with the beam, enabling data-driven modeling of cosmic ray and detector noise. Both simulation and data reconstruction use the Pandora framework Acciarri et al. (2018), which uses a multi-algorithm approach to pattern recognition and event reconstruction in LArTPCs. Pandora creates clusters from wire hits, identifies the neutrino vertex, matches clusters across wire planes to form 3D objects, classifies reconstructed particles as tracks or showers, and establishes particle hierarchies by identifying parent-daughter relationships among particles.
@@
Before performing a dedicated selection for $K^{+}$ production signal events, a pre-selection is applied as follows: (1) events must pass a CC inclusive filter which isolates CC neutrino interactions by identifying the presence of an outgoing muon from the vertex of a neutrino interaction Abratenko et al. (2019), (2) all reconstructed tracks in an event must originate within a 2.4 m × 2.1 m × 9.8 m fiducial volume centered inside the MicroBooNE TPC, and (3) the daughter track (defined as a track connected to the endpoint of one primary signal track candidate) should stop at least 5 cm away from any of the physical edges of the TPC to ensure containment.
@@
For the $K^{+}$ event selection, we implemented a boosted decision tree (BDT) classifier Hoecker et al. (2009). The BDT classifier inputs include reconstructed variables such as $\chi^{2}$ values which compare the expected $dE/dx$ for kaon ($\chi^{2}_{k}$), proton ($\chi^{2}_{p}$), pion ($\chi^{2}_{\pi}$), and muon ($\chi^{2}_{\mu}$) hypotheses with the observed data for each wire plane, the combined 3-wire-plane $\chi^{2}$, the daughter track length, and the log-likelihood ratio particle identification Abratenko et al. (2021). We trained the BDT classifier to identify true $K^{+}$ that decay to $\mu^{+}\nu_{\mu}$ or $\pi^{+}\pi^{0}$, using Monte Carlo (MC) simulation samples of reconstructed neutrino interactions that meet the preselection criteria described above. To train the BDT classifier, a background sample corresponding to $3.67\times 10^{20}$ POT (equivalent to 10% of the MicroBooNE BNB simulation) was used. For the simulation of the signal, a dedicated high-statistics sample of single and associated $K^{+}$ production was used. We optimized the BDT classifier score selection criteria by maximizing the product of efficiency and purity. Figure 2 shows the BDT classifier score distribution, with the orange arrow at 0.41 representing the selected BDT score cut. Events with a score higher than the BDT selection are kept as $K^{+}$ candidates, and 10 such $K^{+}$ candidates are identified in the data. The total number of background events predicted from neutrino-induced interactions, cosmic-rays recorded during neutrino-triggered beam-spills, and off-beam data is 2.21 events. The simulation predicts a selection efficiency of 3.95%, and an associated selection purity of 79%. The primary background category for $K^{+}$ candidates consists of protons generated at the neutrino vertex that interact with an argon nucleus, producing secondary particles at the proton endpoint that could mimic a $K^{+}$ decay product. This background category comprises 2.08 out of 2.21 predicted background events.
@@
To evaluate the accuracy of the background modeling, two sideband regions were defined using a 2-dimensional distribution of the BDT classifier score versus daughter track length. A far sideband region is defined to include tracks with low BDT classifier scores (BDT << 0) and track lengths less than 40 cm. A near sideband region selects tracks with BDT classifier scores close to the signal region (BDT << 0.41) and across all track length ranges, but excludes tracks in the far sideband region. The data/MC comparisons of the near and far sidebands show agreement for reconstructed kinematic variables such as $K^{+}$ candidate track length, $\phi$ (azimuthal angle around the beam direction), and $\theta$ (angle of the track with respect to the beam direction). For the $K^{+}$ candidate track length data/MC comparison, $\chi^{2}/\text{ndf}$ values (including statistical uncertainties) of 16.87/24 and 17.93/24 were obtained for the near and far sidebands. The agreement with the data suggests that the simulation is able to adequately model the background in the cross-section extraction. More information about the sideband comparisons is provided in the Supplemental Material sup .
@@
The systematic uncertainties in the measured cross section are evaluated by reweighting and generating simulated events to account for each source of uncertainty Abratenko et al. (2022a). For the detector-related uncertainties, a dedicated simulation is used to model the variation in detector responses Abratenko et al. (2022b); Calcutt et al. (2021). In contrast, uncertainties from neutrino interactions and BNB flux modeling are treated with a reweighting technique. To evaluate the uncertainty associated with kaon reinteractions within the detector, we implement a methodology similar to that presented by MINERvA Marshall et al. (2016). For each selected $K^{+}$ candidate track, a weight $W_{\text{inel}}$ is calculated using the following expression to estimate the kaon inelastic reinteraction uncertainty:
@@
where $\rho$ is the density of the material, $x$ is the track length, $\sigma^{\text{tot}}_{\text{geant}}$ and $\sigma^{\text{inel}}_{\text{geant}}$ are the $K^{+}$ total and inelastic GEANT4 cross sections, and $\sigma^{\text{tot}}_{\text{var}}$ and $\sigma^{\text{inel}}_{\text{var}}$ are the $K^{+}$ total and inelastic cross sections estimated by applying a variation of $\pm$40% to the $\sigma^{\text{inel}}_{\text{geant}}$. The 40% variation is motivated by observed disagreement between experimental measurements of $K^{+}-\text{Ca}$ interactions Friedman et al. (1997) and the predictions from the GEANT4 interaction model shown in the supplemental material sup .
The leading source of systematic uncertainty comes from the detector response (30.3%), with a major contribution from the modeling of ion recombination of electrons with argon. This analysis relies on identifying stopping $K^{+}$ and $\mu^{+}$ ($\pi^{+}$) through their Bragg peaks, which are characterized by a sharp rise in energy deposition near the end of the particle track. These effects occur in a highly ionizing regime, where ion recombination effects are most pronounced. As the charge profile is impacted by recombination, any mismodeling can distort the Bragg peak shape and affect particle identification. The modeling of neutrino interactions, the BNB neutrino flux, and particle reinteractions in the detector, including kaon inelastic interactions, lead to systematic uncertainties of 3.7%, 11.7%, and 10.5% respectively. The effect of statistics in the estimation of the selection efficiency (6.1%), the number of target nuclei $N_{\text{target}}$ (1%), MC statistical uncertainties on the background estimation (7.4%), and POT exposure (2%) are the remaining sources of uncertainty. The overall magnitude of the uncertainty (statistical and systematic) on our measurement is 54.1%, where the major contribution is due to the statistical uncertainty, with a reported value of 40.6%. Table 1 presents the measured cross section of the $\nu_{\mu}+\text{Ar}\rightarrow\mu^{-}+K^{+}+X$ interactions and the predictions obtained from various neutrino event generators with the NUISANCE framework Stowell et al. (2017).
@@
In conclusion, this letter presents the first measurement of the $\nu_{\mu}+\text{Ar}\rightarrow\mu^{-}+K^{+}+X$ cross section on argon, using the MicroBooNE dataset collected from the BNB flux. The results reported here represent a significant step forward for future studies aiming to refine background estimates for proton decay searches, particularly within the framework of GUTs and SUSY. To advance this analysis in future LArTPC measurements, it is necessary to increase statistics, reduce large detector systematics (especially those related to the recombination of electrons with argon nuclei), refine reconstruction algorithms for short tracks, and improve the identification of kaon tracks that undergo interactions. Experiments such as SBND Acciarri et al. (2025), ICARUS Amerio et al. (2004), and the DUNE near detector will allow precision measurements of neutrino-induced $K^{+}$ production on argon, including differential cross-section measurements for the first time ever. Furthermore, the uncertainty related to kaon reinteractions on argon nuclei could be refined by using data from LArIAT Gramellini (2018), and the DUNE prototypes at CERN Abud et al. (2024). More statistics will come from analyzing MicroBooNE’s full dataset with a total of 1.2 $\times 10^{21}$ POT. The future DUNE near detector data could also provide more accurate measurements of CC neutrino-induced $K^{+}$ production on argon nuclei enabling improved comparisons among the neutrino generator models. These high-precision measurements will enhance our understanding of rare neutrino interactions and final state interaction modeling, and will improve the background predictions for future nucleon decay searches at DUNE, which currently rely on model predictions.

### v4

@@ Abstract
Within the Standard Model (SM), the stability of protons is a consequence of the global symmetry that conserves baryon number. However, various Grand Unified Theories (GUTs) predicting baryon number violation through nucleon decay suggest different decay channels, with a preference for $p\rightarrow e^{+}\pi^{0}$, with lifetimes ranging from $10^{34}$ to $10^{39}$ years PhysRevLett.32.438. Similarly, the Supersymmetric Grand Unified Theory (SUSY GUT) favors protons decaying via $p\rightarrow\nu K^{+}$, with lifetimes of $10^{32}$ to $10^{35}$ years Hisano:1992jj.
Upcoming large long-baseline neutrino oscillation experiments, such as the Deep Underground Neutrino Experiment (DUNE) dunetdr and Hyper-Kamiokande protocollaboration2018hyperkamiokandedesignreport, have the sensitivity to search for proton decay in the range of $10^{32}$ to $10^{39}$ years. Cherenkov detector experiments like Super-Kamiokande PhysRevD.90.072005; PhysRevD.95.012004; PhysRevD.102.112011 and Hyper-Kamiokande are particularly suited for $p\rightarrow e^{+}\pi^{0}$ decay searches. However, for water Cherenkov detectors, the $p\rightarrow\nu K^{+}$ decay channel needs to be indirectly detected since the momentum of the produced $K^{+}$ (339 MeV/c) is below the Cherenkov detection threshold in water (560 MeV/c). Furthermore, backgrounds from atmospheric neutrino interactions, such as the $\nu_{\mu}n\rightarrow\mu^{-}K^{+}n$ and $\nu p\rightarrow\nu K^{+}\Lambda$ channels, are challenging to distinguish in water Cherenkov detectors.
@@
The first measurements of charged-current (CC) and neutral-current (NC) kaon production, reported with the Argonne National Laboratory PhysRevLett.33.1446 and GARGAMELLE bubble chambers osti_4109186 in 1974 and 1975, observed $\mathcal{O}$(10) events. Later, in the 1980’s, the Brookhaven National Laboratory bubble chamber reported $\mathcal{O}$(10) CC and NC events with strange particle production in neutrino-hydrogen and neutrino-deuterium interactions Jensen:1964zz; PhysRevD.24.2779, and the Fermilab bubble chamber reported $\mathcal{O}$(10) events and a measured cross section for CC neutrino interactions $\nu_{\mu}+n\xrightarrow{}\mu^{-}+\Lambda+K^{+}$ on a deuterium target for energy ranges between 10 GeV and 250 GeV PhysRevD.28.2129. Most recently, the MINERvA experiment measured $K^{+}$ production in NC and CC modes on a hydrocarbon target, reporting $\mathcal{O}$(100) and $\mathcal{O}$(1000) events respectively PhysRevD.94.012002; PhysRevLett.119.011802; Marshall:2016mar. In addition, MINERvA reported the first evidence of coherent kaon production PhysRevLett.117.061802; Marshall:2016mar.
@@
The MicroBooNE detector Fleming:2012gvl; MicroBooNE:2016pwy consists of an 85 tonne LArTPC and a photon detection system comprised of 32 photomultiplier tubes (PMTs). The detector identifies neutrino interactions by detecting ionization and scintillation light produced by charged particles moving through the detector. Ionization charge is recorded across three wire planes (two induction, one collection) oriented at different angles (with induction planes at $\pm 60^{\circ}$ relative to the vertical collection plane), and 3 mm wire-spacing. This configuration allows the generation of mm-scale resolution, three-dimensional images of neutrino interactions. Scintillation light collected by the PMTs provides ns-scale timing information, essential for identifying neutrino interactions that are synchronized with the BNB, and for rejecting cosmic-ray backgrounds.
@@
In MicroBooNE, the LArSoft framework Snider:2017wjd provides tools for simulating the neutrino flux, neutrino interactions, particle propagation, and detector response including ionization and scintillation processes as well as reconstruction of neutrino events. The BNB neutrino flux through the MicroBooNE detector is simulated using the MiniBooNE flux model PhysRevD.79.072002 adapted for MicroBooNE’s location along the beamline. GENIE v3.0.6 (G18_10a_02_11a) Andreopoulos:2009rq; PhysRevD.104.072009 is the event generator used to simulate neutrino-argon interactions, including both single and associated kaon production. In GENIE, associated kaons are produced via individual resonances simulated by the Rein-Sehgal model REIN198179. Kaons produced via deep inelastic scattering are simulated by the Koba-Nielsen-Olesen (KNO) scaling parameterization KOBA1972317. For higher energies (hadronic invariant mass W >> 2.3 GeV), the AGKY model Yang_2009 and PYTHIA6 Sj_strand_2006 are used. The hadronization models are tuned to match strange particle production ($\Lambda$ and $K^{0}_{S}$) observed in experiments such as the Big European Bubble Chamber Jones:1992bm; ALLASIA19831; WA59:1991cop; BOSETTI198229 and the Fermilab bubble chamber PhysRevD.34.1251; PhysRevD.50.6691. The single kaon production model skmodel is implemented in GENIE, generating events in the channels $\nu_{l}+n\rightarrow l^{-}+n+K^{+}$, $\nu_{l}+p\rightarrow l^{-}+p+K^{+}$, and $\nu_{l}+n\rightarrow l^{-}+p+K^{0}$. The region of validity assumed for the model is for $E_{\nu}\leq$ 2 GeV skmodel; skmodel2, while the GENIE authors note that the model can be applied to higher energies as well geniesingle.
@@
Particle trajectories and interactions within the detector are simulated using GEANT4 v4_10_3_p03c using the QGSP_BERT hadron physics list AGOSTINELLI2003250; 1610988; ALLISON2016186, which models particle propagation in the detector. Simulated neutrino interactions are overlayed with data events collected using an unbiased trigger that operates in anti-coincidence with the beam, enabling data-driven modeling of cosmic ray and detector noise. Both simulation and data reconstruction use the Pandora framework pandora, which uses a multi-algorithm approach to pattern recognition and event reconstruction in LArTPCs. Pandora creates clusters from wire hits, identifies the neutrino vertex, matches clusters across wire planes to form 3D objects, classifies reconstructed particles as tracks or showers, and establishes particle hierarchies by identifying parent-daughter relationships among particles.
@@
Before performing a dedicated selection for $K^{+}$ production signal events, a pre-selection is applied as follows: (1) events must pass a CC inclusive filter which isolates CC neutrino interactions by identifying the presence of an outgoing muon from the vertex of a neutrino interaction MicroBooNE:2019nio, (2) all reconstructed tracks in an event must originate within a 2.4 m × 2.1 m × 9.8 m fiducial volume centered inside the MicroBooNE TPC, and (3) the daughter track (defined as a track connected to the endpoint of one primary signal track candidate) should stop at least 5 cm away from any of the physical edges of the TPC to ensure containment.
@@
For the $K^{+}$ event selection, we implemented a boosted decision tree (BDT) classifier tmvaroot. The BDT classifier inputs include reconstructed variables such as $\chi^{2}$ values which compare the expected $dE/dx$ for kaon ($\chi^{2}_{k}$), proton ($\chi^{2}_{p}$), pion ($\chi^{2}_{\pi}$), and muon ($\chi^{2}_{\mu}$) hypotheses with the observed data for each wire plane, the combined 3-wire-plane $\chi^{2}$, the daughter track length, and the log-likelihood ratio particle identification nicpaper2021. We trained the BDT classifier to identify true $K^{+}$ that decay to $\mu^{+}\nu_{\mu}$ or $\pi^{+}\pi^{0}$, using Monte Carlo (MC) simulation samples of reconstructed neutrino interactions that meet the preselection criteria described above. To train the BDT classifier, a background sample corresponding to $3.67\times 10^{20}$ POT (equivalent to 10% of the MicroBooNE BNB simulation) was used. For the simulation of the signal, a dedicated high-statistics sample of single and associated $K^{+}$ production was used. We optimized the BDT classifier score selection criteria by maximizing the product of efficiency and purity. Figure 2 shows the BDT classifier score distribution, with the orange arrow at 0.41 representing the selected BDT score cut. Events with a score higher than the BDT selection are kept as $K^{+}$ candidates, and 10 such $K^{+}$ candidates are identified in the data. The total number of background events predicted from neutrino-induced interactions, cosmic-rays recorded during neutrino-triggered beam-spills, and off-beam data is 2.21 events. The simulation predicts a selection efficiency of 3.95%, and an associated selection purity of 79%. The primary background category for $K^{+}$ candidates consists of protons generated at the neutrino vertex that interact with an argon nucleus, producing secondary particles at the proton endpoint that could mimic a $K^{+}$ decay product. This background category comprises 2.08 out of 2.21 predicted background events.
@@
To evaluate the accuracy of the background modeling, two sideband regions were defined using a 2-dimensional distribution of the BDT classifier score versus daughter track length. A far sideband region is defined to include tracks with low BDT classifier scores (BDT << 0) and track lengths less than 40 cm. A near sideband region selects tracks with BDT classifier scores close to the signal region (BDT << 0.41) and across all track length ranges, but excludes tracks in the far sideband region. The data/MC comparisons of the near and far sidebands show agreement for reconstructed kinematic variables such as $K^{+}$ candidate track length, $\phi$ (azimuthal angle around the beam direction), and $\theta$ (angle of the track with respect to the beam direction). For the $K^{+}$ candidate track length data/MC comparison, $\chi^{2}/\text{ndf}$ values (including statistical uncertainties) of 16.87/24 and 17.93/24 were obtained for the near and far sidebands. The agreement with the data suggests that the simulation is able to adequately model the background in the cross-section extraction. More information about the sideband comparisons is provided in the Supplemental Material suppmat.
@@
The systematic uncertainties in the measured cross section are evaluated by reweighting and generating simulated events to account for each source of uncertainty MicroBooNE:2021ccs. For the detector-related uncertainties, a dedicated simulation is used to model the variation in detector responses Abratenko_2022; Calcutt_2021. In contrast, uncertainties from neutrino interactions and BNB flux modeling are treated with a reweighting technique. To evaluate the uncertainty associated with kaon reinteractions within the detector, we implement a methodology similar to that presented by MINERvA PhysRevD.94.012002. For each selected $K^{+}$ candidate track, a weight $W_{\text{inel}}$ is calculated using the following expression to estimate the kaon inelastic reinteraction uncertainty:
@@
where $\rho$ is the density of the material, $x$ is the track length, $\sigma^{\text{tot}}_{\text{geant}}$ and $\sigma^{\text{inel}}_{\text{geant}}$ are the $K^{+}$ total and inelastic GEANT4 cross sections, and $\sigma^{\text{tot}}_{\text{var}}$ and $\sigma^{\text{inel}}_{\text{var}}$ are the $K^{+}$ total and inelastic cross sections estimated by applying a variation of $\pm$40% to the $\sigma^{\text{inel}}_{\text{geant}}$. The 40% variation is motivated by observed disagreement between experimental measurements of $K^{+}-\text{Ca}$ interactions PhysRevC.55.1304 and the predictions from the GEANT4 interaction model shown in the supplemental material suppmat.
[…]
@@
In conclusion, this letter presents the first measurement of the $\nu_{\mu}+\text{Ar}\rightarrow\mu^{-}+K^{+}+X$ cross section on argon, using the MicroBooNE dataset collected from the BNB flux. The results reported here represent a significant step forward for future studies aiming to refine background estimates for proton decay searches, particularly within the framework of GUTs and SUSY. To advance this analysis in future LArTPC measurements, it is necessary to increase statistics, reduce large detector systematics (especially those related to the recombination of electrons with argon nuclei), refine reconstruction algorithms for short tracks, and improve the identification of kaon tracks that undergo interactions. Experiments such as SBND sbndcollaboration2025shortbaselineneardetectorfermilab, ICARUS AMERIO2004329, and the DUNE near detector will allow precision measurements of neutrino-induced $K^{+}$ production on argon, including differential cross-section measurements for the first time ever. Furthermore, the uncertainty related to kaon reinteractions on argon nuclei could be refined by using data from LArIAT osti_1489387, and the DUNE prototypes at CERN PhysRevD.110.092011. More statistics will come from analyzing MicroBooNE’s full dataset with a total of 1.2 $\times 10^{21}$ POT. The future DUNE near detector data could also provide more accurate measurements of CC neutrino-induced $K^{+}$ production on argon nuclei enabling improved comparisons among the neutrino generator models. These high-precision measurements will enhance our understanding of rare neutrino interactions and final state interaction modeling, and will improve the background predictions for future nucleon decay searches at DUNE, which currently rely on model predictions.